# This work was supported by project from DOE-CESER

### v1

@@ V-B Capacitor Bank
Capacitor banks help with providing the reactive power support to the grid to increase the maximum power transfer capability of the system. Generally, capacitor banks are a group of capacitor units connected in series and parallel circuits to achieve desired voltage and VAR rating respectively. When a capacitor unit fails, it causes unbalance and results in tripping of capacitor bank. Without protection trip, the capacitor units in the bank results in a cascading failure [74].
Excess heat accumulation is one of the common causes of a capacitor bank failure. Therefore, the internal temperature is used as the physical quantity to identify failure of a capacitor bank. Furthermore, this internal temperature is function of both ambient temperature and the radiant heat from the nearby equipment [75]. Thus ambient temperature not only helps with monitoring of capacitor bank failure but also effects the life of capacitor bank [75, 76].
Another indirect reason for capacitor bank failures is due to load growth on an extremely hot day. The temperature rise of a capacitor is a function of the total internal watts dissipated by the windings. Initially, the watts are due to fundamental current (squared) and equivalent series resistance of the windings (ESR). Thus the geometry of the capacitor windings plays a major role.
@@
Existing Methods to Measure the Impacts: Heat impacts the electronic components and likewise effects the capacitor bank failures as well. As shown in Fig. 15 and Fig. 16, the internal temperature of the capacitor bank is not only impacted by over voltage (lightly loaded - hot/cold days depending climate zone) and over current (heavily loaded - hot/cold days depending climate zone) but also by the geometric dimensions of the capacitor winding [77][79]. The overall operating temperature is a function of internal and ambient temperature. Thus depending on the internal temperature, the limits on ambient temperature are strictly imposed. It is to be noted that the ambient temperature referred here is not the room’s air but it refers to the temperature within the capacitor bank’s enclosure i.e., the air that is surrounding the capacitor bank. Generally, this can be 20°C above the ambient room’s air temperature. Therefore, monitoring the enclosure temperature during extreme heat helps to identify impacts on the capacitor bank. From Fig. 17, for every 7°C above the rated operating temperature, the capacitor life is reduced by 50% [77].
Apart of temperature monitoring, exposure to sharp temperature variations, transient over voltages, aging and manufacturing defects can cause internal failures of capacitor elements. A new method using indicating quantity Super imposed Reactance (SR), is presented in [80].
Existing Methods to Mitigate the Impacts: Capacitors that are often operated under uncertain loading conditions when placed in a climate controlled environment helps mitigate failure of capacitor banks. A study of capacitor internal heat generation, reported in IEEE Transactions on Industry Applications [81] revealed that for a given capacitance and winding area, short windings of larger diameter generated far less internal temperature rise than tall windings. The article compares four capacitors with equal $\mu$F, voltage rating, film thickness and thickness of metallization thickness. The only difference is the height of the four capacitors. The conclusion was that the optimized winding construction for capacitors is a short coil which results in a relatively larger diameter. Although it may be less expensive to produce capacitors with tall windings, they do not handle heat as well as short windings.
Use a reactor in series with capacitors whenever harmonics are present on the electrical system to which capacitors will be applied. Use 7% detuning reactor (tunes capacitor at $3.78^{th}$ harmonic) to protect power factor capacitors [82]. This reduces heat as it reduces over current in the line.
@@ V-C Batteries
There are various types of batteries depending on their construction and the electrolyte material used in the batteries. In this work, we do not extensively focus on the electrolyte material but rather discuss about how temperature impact the efficiency of the batteries and how the impact varies between different types of batteries. According to EIA, in 2022, global investments in battery energy storage exceeded 20 billion US dollars and 65% of it is for grid-scale deployment. It was projected that battery energy storage investments to exceed 35 billion US dollars in 2023 [83].
Based on cost and energy density considerations, lithium iron phosphate batteries, a subset of lithium-ion batteries, are still the preferred choice for grid-scale storage. For applications with limited space and home energy storage, energy-dense chemistries for lithium-ion batteries, such as nickel cobalt aluminium (NCA) and nickel manganese cobalt (NMC) are popular. However, apart from lithium-ion batteries, flow batteries with low performance-degradation are also under development/research for technology breakthrough energy storage needs with limited space. In 2022, world’s largest vanadium redox flow battery was comissioned in China with capacity of 100 MW and energy volume of 400 MWh [83]. At the end of 2022, the entire U.S. grid had more than 1.1 million MW of capacity [84].
Impact of Climate Change: Storage solutions for grid are either placed in a climate control indoor setup or outdoors with appropriate cooling setup [85]. Primarily, high temperatures cause more depletion of energy in a battery. Thus, extreme temperature in combination with any possible point of failure in cooling system can result in a battery performance that may not be sufficient to support the grid functions. [86] reports the failure of grid-scale battery packs in Australia due to high temperature and failure of cooling system.
Another popular cause for storage system failure is an indirect one where the wildfire cause transmission faults which indirectly impact the PV resources [87]. The most leading cause for BESS failures is fire based [88]. Fig. 18 shows the increase in BESS related events increase in recent years. This is not surprising since the BESS had a tremendous growth in recent years and from Fig. 19, we can observe that the BESS events are often observed when their age between 0 and 3 years. Another useful statistic is shown to be that as the system age increases, the number of events reduces [88]. More than 50% of BESS failures occur within the initial two years of operation. 5-50 MWh BESS account for over half of total failure events globally [89]. 48% of failures have been linked to solar-plus-storage projects [89]. In addition, some locations subjected to rapid temperature variations such as in the mountains can experience dewing leading to damage within the ESS located outdoors if not well-controlled [90].
@@
Existing Methods to Measure the Impacts: The existing method to measure the impact of extreme temperatures on BESS are through battery management systems (BMS). BMS monitors the temperature of different cells in a battery bank, coolant intake/output temperature, state of the battery, etc. [91] presents how different temperatures impact the performance parameters of a battery with a testbed. Extreme and unexpected change in temperatures can cause sub zero temperatures which are shown to impact the capacity loss [92].
Existing Methods to Mitigate the Impacts: Some ways to mitigate the impacts are as follows [89]. 1) Ensuring sufficient spacing between battery modules is critical as radiant heat impact from adjacent equipment and their failure can cause cascading failures, 2) Carefully selecting the right type of battery for BESS depending on the local climate, 3) Creating spacing standards for BESS units, and 4) Involving OEMs throughout the entire BESS project lifecycle.
Furthermore, the other solution to mitigate the impacts is through a reliable BMS and redundant climate control options to maintain the temperatures of the BESS within the permissible limits. However, redundant systems are more expensive to invest and maintain especially when it involves BESS. There has been work on early prediction of battery life [93]. [94] describes a method employed by a commercially available battery monitoring system which, through impedance and other measurements, provides the user with a viable and predictive approach to improving overall system reliability. [95] presents a prediction methodology to identify faults by using redundant sensor.
@@ VI-A Transformer
Transformers are critical infrastructure for transmission and distribution circuits of power grids. Transformers used in power domain can be segregated as 1) power transformers, 2) measurement transformers (current and potential transformers), and 3) distribution transformers.
@@
    Power Transformers: Power transformers facilitate energy transfer between high voltage transmission system and primary distribution system. These transformers are generally large and are located in substations. Based on operational voltage level, these transformers can be segregated into three types, namely 1) small power transformer (up to 7 MVA), 2) medium power transformer (50 – 100 MVA), and 3) large power transformer (more than 100 MVA).
@@
    Measurement transformers: Measurement transformers such as current and potential transformers are used to measure the current and voltage of the power line.
@@
    Distribution transformers: Distribution transformers facilitate energy transformer at very low voltage when compared to power transformers. More specifically, 33kV for industrial applications and 440-220V for residential applications.
@@
Impact of Climate Change: The impact of extreme temperatures on transformers can be understood by studying the direct and indirect impacts of the ambient temperature. Direct impact involves the ambient temperature’s impact on transformer components. Whereas indirect impact involves the ambient temperature’s impact on electrical loading of the transformer which in turn causes overheating. For example, the winding and oil inside the transformer gets hotter as the loading and losses increase. Through conduction, the heat that arises in the winding impacts the transformer oil that is used as a coolant. Excessive heat will damage the insulation resulting in transformer breakdown.
@@
Power transformer is one of the vital equipment in the substation. Power transformer’s operation is susceptible to temperature changes mainly two ways. First is due to changes in ambient temperature causes temperature of coolant oil to change resulting in transformer ageing and sometimes even a breakdown [96]. Second is due to combined effect of direct impact and indirect impact of ambient temperature on the transformer [96]. Heat can cause anomaly in some part of the transformer which can increase ageing of transformer [97]. According to the IEC standard, power transformers are used at an ambient temperature of 20°C [98] (30°C IEEE standard [99]). However, environmental temperature differences in design and real conditions make a transformer very risky to operate to reach its maximum capacity [100]. [101] shows that ambient temperature changes also impact breakdown voltage of the transformer. For example, the allowable temperature rise of coil is 65°C. Based on Class A insulation for winding, maximum operating temperature is 105°C. When the ambient temperature is 40°C, the allowable temperature rises to 65°C (105°C - 40°C = 65°C). As the temperature of transformer is generally 10°C lower than that of winding, the allowable temperature rise of transformer oil is 55°C. To prevent oil aging, the temperature rise of upper oil surface shall not exceed 45°C. No matter how the ambient air changes, the safe operation of the transformer within the specified service life can be guaranteed only if the temperature rise does not exceed the allowable value of 45°C.
@@
Instrument/measurement transformers play a vital role for measurement and protection in power systems. It is reported that the accuracy of instrument transformers is impacted by the ambient temperature [102, 103]. For example, the inductive current transformer does not fulfill the accuracy class requirements when the temperature is lower the conventional ambient temperature of 23°C but higher than the minimum working temperature specified by the standards [104]. Industry also reports the derating nature of the current transformers due to increase in ambient temperature [105].
@@
Data from real-world (Chicago, Illinois) is used to study the effect of ambient temperature on distribution transformers, it is shown that the overall temperature of the transformer is affected by ambient temperature and load level [106]. Furthermore, the ratings of the distribution transformer need to be designed based on the highest temperature recorded and altitude of the site [107]. [108] also shows how the heat generation within the padmount transformer leads to failures.
@@
Existing Methods to Measure the Impacts: The existing methods measure the impacts of extreme temperatures on transformers by developing techniques to quantify the ageing of transformer, and failure prediction for insulation breakdown. [109] presents a methodology to estimate the ageing of transformer by predicting transformer temperature rise. Industry also focuses on monitoring the temperature and moisture of the winding insulation to track the aging rate using dynamic rate tracking [110]. Furthermore, for every 6°C increase in winding temperature above the rated maximum of 110°C, the transformer aging rate roughly doubles [110]. There are also thermal imaging solutions that are used to monitor the temperature of the transformer which is further used to calculate the internal winding and oil temperature to determine the ageing impact or potential failure [111, 112, 113]. [114, 115] developed a prediction model that uses weather data to correlate the distribution transformer failure. Specifically, [115] provides three metrics that can be used with the help of a closed form expression obtained in the work to study the impacts on transformer life expectancy for any region.
@@
Existing Methods to Mitigate the Impacts: The various methods to prevent overheating of a transformer includes load sizing that is supported by the transformer, auxiliary cooling systems, adequate ventilation (placement proximity of transformer and other equipment), insulation inspection, oil quality checks, and regular maintenance. Out of the various solutions, auxiliary cooling systems and adequate ventilation plays a crucial role when it comes to extreme temperature impacts on the transformers. Please note that the auxiliary cooling systems are additional systems that are not part of the transformer’s design i.e., excludes cooling methods like AN, AF, ONAN, ONAF, etc. [116]. There have been mobile transformers that are used to mitigate the power grid impacts due to transformer failure in extreme temperatures [117], this enabled to reduce the down time significantly. Another solution is to provide mobile auxiliary cooling services, for example, one substation in the eastern United States was facing 50,000 gallons of overheated transformer oil due to an overloaded transformer where the typical 208V/240V cooling application is not sufficient. In this case, portable trailer-mounted motor cooling is used to cool the transformer [118]. Another well known substation solution is thermal imagery techniques to monitor the temperatures of the transformer and its adjacent equipment which can also radiate heat increasing transformer’s temperature [111, 112, 113].
@@ VI-B Circuit Breaker
Circuit breakers are essential grid equipment whose role is to protect the system and isolate disturbances or faults in the power grid. Circuit breakers are primary grid equipment that also carry power and will have internal device heat during heavy loading conditions on hot days. The insulating medium in circuit breakers change with the system voltage and has a role to play in reliable operation of a circuit breakers. High voltage circuit breakers usually have an insulating medium: SF6 gas, nitrogen, mix of these gases, vacuum circuit breakers are the common high and medium voltage circuit breakers. Medium and low voltage circuit breakers are usually air insulated circuit breakers. While reference [119] highlights that extreme cold conditions will degrade the performance of SF6 circuit breakers, reference [120] concludes that extreme heat and temperature rise of SF6 has little impact on the SF6 gas performance and its application in SF6 insulated power equipment. There are other studies in literature that have studied some gas mixtures like SF6 - CO2 as insulating media in circuit breakers but do not discuss the performance under extreme temperatures [121].
@@
Impact of Climate Change: Medium voltage and low voltage circuit breakers that are mainly for industrial and residential protection, have significant impact on their performance and behavior during extreme heat and cold. Reference [122] provides details of de-rating and re-rating curves for low voltage residential circuit breakers up to 63 A ratings. These clearly indicate that extreme heat and hot ambient air conditions will negatively impact the performance of the circuit breakers and can lead to tripping of the breakers much earlier than intended leading to loss of load and often nuisance tripping. Aggregated impact of such behavior can cause significant impact on the aggregated load in the system leading to unwanted cascaded actions on the distribution system level. Reference [123] discusses the reliability concerns of other low voltage high power circuit breakers in extreme weather conditions. Evaluating the temperatures experienced during transportation and storage is an important consideration for circuit breakers, and needs to be documented to ensure the grid operators and asset owners understand the expected life and reliability vulnerability of the equipment that is being commissioned.
@@
Reference [124] is a report by NERC that discusses examples of circuit breaker failures due to extreme cold that are reported. These occurrences are important to consider in various regions of the country as the unstable polar vortex phenomenon is happening with increasing frequency and reaching regions of the country that normally did not expect cold days of the year.
@@
Existing Methods to Measure the Impacts: Reference [125] Provides details of some methods for monitoring SF6 gas conditions in SF6 circuit breakers. Reference [126] provides a circuit breaker monitoring system that asset owners can utilize to monitor various operating conditions of the circuit breakers and can itself operate in temperatures ranging from -40°C to +65°C which means that it might need additional thermal insulation to operate in extreme temperature exposures like during polar vortex or extreme hot days inside an enclosure. Other details specifications of this monitoring system can be found in reference [127].
@@
Existing Methods to Mitigate the Impacts: While extreme heat related failures can be managed to some extent by derating that compensates the temperature rise casued due to loading of the equipment, extreme cold conditions is difficult to be managed for outdoor equipment. Reference [7] provides links to other learning and reported material associated with winter preparedness for circuit breakers and other critical grid equipment. Winterising the equipment that depend on fluids is important and oftentimes high voltage circuit breakers have some media around the contacts to extinguish the arc and break the circuit [124].

### v2

@@ V-B Capacitor Bank
Capacitor banks help with providing reactive power support to the grid to increase the maximum power transfer capability of the system. Generally, capacitor banks are a group of capacitor units connected in series and parallel circuits to achieve desired voltage and VAR rating, respectively. When a capacitor unit fails, it causes unbalance and results in tripping of capacitor bank. Without a protection trip, the capacitor units in the bank would experience a cascading failure [73].
Excess heat accumulation is one of the common causes of a capacitor bank failure. Therefore, the internal temperature is used as the physical quantity to identify failure of a capacitor bank. This internal temperature is function of both ambient temperature and the radiant heat from the nearby equipment [74]. Thus, ambient temperature not only helps with monitoring of capacitor bank failure but also effects the life of capacitor bank [74, 75].
Another indirect reason for capacitor bank failures is because of load growth on an extremely hot day. The temperature rise of a capacitor is a function of the total internal watts dissipated by the windings. Initially, the watts are due to fundamental current (squared) and equivalent series resistance of the windings (ESR). Thus, the geometry of the capacitor windings plays a major role.
@@
Existing Methods to Measure the Impacts: Heat impacts the electronic components and likewise affects the capacitor bank failures as well. As shown in Fig. 15 and Fig. 16, the internal temperature of the capacitor bank is not only impacted by overvoltage (lightly loaded – hot/cold days depending climate zone) and over current (heavily loaded – hot/cold days depending climate zone) but also by the geometric dimensions of the capacitor winding [76][78]. The overall operating temperature is a function of internal and ambient temperature. Thus depending on the internal temperature, the limits on ambient temperature are strictly imposed. Note that the ambient temperature referred here is not the room’s air but the temperature within the capacitor bank’s enclosure, i.e., the air that is surrounding the capacitor bank. Generally, this can be 20°C above the ambient room’s air temperature. Therefore, monitoring the enclosure temperature during extreme heat helps to identify impacts on the capacitor bank. From Fig. 17, for every 7°C above the rated operating temperature, the capacitor life is reduced by 50% [76].
Apart from temperature monitoring, exposure to sharp temperature variations, and transient overvoltages, aging and manufacturing defects can cause internal failures of capacitor elements. A new method using indicating quantity superimposed reactance (SR) is presented in [79].
Existing Methods to Mitigate the Impacts: Capacitors that are often operated under uncertain loading conditions can be placed in a climate-controlled environment to help mitigate the risk of failure. A study of capacitor internal heat generation, reported in IEEE Transactions on Industry Applications [80], revealed that for a given capacitance and winding area, short windings of larger diameter generated far less internal temperature rise than tall windings. The article compares four capacitors with equal $\mu$F, voltage rating, film thickness and thickness of metallization. The only difference is the height of the four capacitors. The conclusion was that the optimized winding construction for capacitors is a short coil, which results in a relatively larger diameter. Although it may be less expensive to produce capacitors with tall windings, they do not handle heat as well as short windings.
Further, it is recommended to use a reactor in series with capacitors whenever harmonics are present on the electrical system to which capacitors will be applied. Use a 7% detuning reactor (tunes capacitor at $3.78^{\textnormal{th}}$ harmonic) to protect power factor capacitors [81]. This reduces heat by reducing over current in the line.
@@ V-C Batteries
Batteries having varying constructions, anode and cathode materials, and electrolyte compositions. This study does not focus on the electrolyte material but rather discusses how temperature affects battery efficiency and how the impact varies between different types of batteries. According to EIA, in 2022, global investments in battery energy storage exceeded 20 billion US dollars, and 65% of it is for grid-scale deployment. Battery energy storage investments are projected to exceed 35 billion US dollars in 2023 [82].
Based on cost and energy density considerations, lithium–iron phosphate batteries, a subset of lithium–ion batteries, are still the preferred choice for grid-scale storage. For applications with limited space and home energy storage, energy-dense chemistries for lithium–ion batteries such as nickel cobalt aluminium (NCA) and nickel manganese cobalt (NMC) are popular. However, flow batteries with low performance degradation are also under development for technology breakthrough energy storage needs with limited space. In 2022, the world’s largest vanadium redox flow battery was commissioned in China with a capacity of 100 MW and energy volume of 400 MWh [82]. At the end of 2022, the entire U.S. grid had more than 1.1 million MW of capacity [83].
Impact of Extreme Temperature Operations : Storage solutions for grid are either placed in a climate-controlled indoor setup or outdoors with an appropriate cooling setup [84]. High temperatures cause more depletion of energy in a battery. Thus, extreme temperature in combination with any possible point of failure in cooling system can result in battery performance that may not be sufficient to support grid functions. Ref. [85] reports the failure of grid-scale battery packs in Australia due to high temperature and failure of cooling systems.
Another common cause for storage system failure is an indirect one: wildfire-caused transmission faults [86]. Fires are the leading cause of BESS failures [87]. Fig. 18 shows the increase in BESS-related events increase in recent years. This is not surprising, because BESS deployment has grown tremendously in recent years. From Fig. 19, we can observe that the BESS events are often observed when their age between 0 and 3 years. Another useful statistic is that as the system age increases, the number of events reduces [87]. More than 50% of BESS failures occur within the initial two years of operation. 5–50 MWh BESS account for over half of total failure events globally [88]. 48% of failures have been linked to solar-plus-storage projects [88]. In addition, some locations subjected to rapid temperature variations such as in the mountains can experience dewing leading to damage within the BESS located outdoors if not well controlled [89].
@@
Existing Methods to Measure the Impacts: The existing method to measure the impact of extreme temperatures on BESS are through battery management systems (BMS). BMS monitors the temperature of different cells in a battery bank, coolant intake and output temperature, state of the battery, etc. Ref. [90] presents how different temperatures affect the performance parameters of a battery with a testbed. Sub-zero temperatures are shown to cause capacity loss [91].
Existing Methods to Mitigate the Impacts: Some ways to mitigate the impacts are as follows [88]. 1) Ensuring sufficient spacing between battery modules is critical as radiant heat impact from adjacent equipment and their failure can cause cascading failures. 2) Carefully selecting the right type of battery for BESS depending on the local climate. 3) Creating spacing standards for BESS units. and 4) Involving OEMs throughout the entire BESS project life cycle.
Furthermore, the other solution to mitigate the impacts is through a reliable BMS and redundant climate control options to maintain the temperatures of the BESS within permissible limits. However, redundant systems are more expensive to acquire and maintain. There has been work on early prediction of battery life [92]. Ref. [93] describes a method employed by a commercially available battery monitoring system which, through impedance and other measurements, provides the user with a viable and predictive approach to improving overall system reliability. [94] presents a prediction methodology to identify faults by using redundant sensors.
@@ VI-A Transformer
Transformers are critical infrastructure for transmission and distribution circuits of power grids. Transformers used in the power domain can be categorized as 1) power transformers, 2) measurement transformers (current and potential transformers), and 3) distribution transformers.
@@
    Power Transformers: Power transformers facilitate energy transfer between the high-voltage transmission system and the primary distribution system. These transformers are generally large and are located in substations. Based on operational voltage level, these transformers can be segregated into three types, namely 1) small power transformer (up to 7 MVA), 2) medium power transformer (50-–100 MVA), and 3) large power transformer (more than 100 MVA).
@@
    Measurement transformers: Measurement transformers, such as current and potential transformers, are used to measure the current and voltage of the power line.
@@
    Distribution transformers: Distribution transformers facilitate energy transfer at very low voltage when compared to power transformers. More specifically, 33 kV for industrial applications and 440–220 V for residential applications.
@@
Impact of Extreme Temperature Operations: The impact of extreme temperatures on transformers can be understood by studying the direct and indirect impacts of the ambient temperature. Direct impact involves the ambient temperature’s impact on transformer components. Whereas indirect impact involves the ambient temperature’s impact on electrical loading of the transformer, which in turn causes overheating. For example, the winding and oil inside the transformer gets hotter as the loading and losses increase. Through conduction, the heat that arises in the winding impacts the transformer oil that is used as a coolant. Excessive heat will damage the insulation, resulting in transformer breakdown.
@@
Power transformers’ operation is susceptible to temperature changes in two main ways. First, changes in ambient temperature can cause the temperature of coolant oil to change, resulting in transformer aging and sometimes even breakdown [95]. Second, the direct impact and indirect impact of ambient temperature can have a combined effect on the transformer [95]. Heat can cause anomalies in some parts of the transformer which, can increase agiing of transformer [96]. According to the IEC standard, power transformers are used at an ambient temperature of 20°C [97] (30°C IEEE standard [98]). However, environmental temperature differences in design and real conditions make a transformer very risky to operate at its maximum capacity [99]. Ref. [100] shows that ambient temperature changes also affect breakdown voltage of the transformer. For example, the allowable temperature rise of a coil is 65°C. Based on Class A insulation for winding, the maximum operating temperature is 105°C. When the ambient temperature is 40°C, the allowable temperature rises to 65°C (105°C - 40°C = 65°C). As the temperature of transformer is generally 10°C lower than that of winding, the allowable temperature rise of transformer oil is 55°C. To prevent oil aging, the temperature rise of upper oil surface shall not exceed 45°C. No matter how the ambient air changes, the safe operation of the transformer within the specified service life can be guaranteed only if the temperature rise does not exceed the allowable value of 45°C.
@@
Instrument/measurement transformers play a vital role for measurement and protection in power systems. The accuracy of instrument transformers is impacted by the ambient temperature [101, 102]. For example, an inductive current transformer does not fulfill the accuracy class requirements when the temperature is lower the conventional ambient temperature of 23°C but higher than the minimum working temperature specified by the standards [103]. Industry also reports the derating nature of the current transformers due to increase in ambient temperature [104].
@@
Data from Chicago, Illinois was used to study the effect of ambient temperature on distribution transformers, and it was shown that the overall temperature of the transformer is affected by ambient temperature and load level [105]. Furthermore, the ratings of the distribution transformer need to be designed based on the highest temperature recorded and altitude of the site [106]. Ref. [107] also shows how the heat generation within the padmount transformer leads to failures.
@@
Existing Methods to Measure the Impacts: Techniques exist to quantify the aging of transformers and predict insulation breakdown. Ref. [108] presents a methodology to estimate the aging of transformers by predicting transformer temperature rise. Industry also focuses on monitoring the temperature and moisture of the winding insulation to track the aging rate using dynamic rate tracking [109]. For every 6°C increase in winding temperature above the rated maximum of 110°C, the transformer aging rate roughly doubles [109]. There are also thermal imaging solutions that are used to monitor the temperature of the transformer, which is further used to calculate the internal winding and oil temperature to determine the aging impact or potential failure [110, 111, 112]. Refs. [113, 114] developed a prediction model that uses weather data to correlate the distribution transformer failure. Specifically, [114] provides three metrics that can be used with the help of a closed form expression obtained in the work to study the impacts on transformer life expectancy for any region.
@@
Existing Methods to Mitigate the Impacts: The various methods to prevent overheating of a transformer includes load sizing that is supported by the transformer, auxiliary cooling systems, adequate ventilation (placement proximity of transformer and other equipment), insulation inspection, oil quality checks, and regular maintenance. Out of the various solutions, auxiliary cooling systems and adequate ventilation play a crucial role in preventing extreme temperature impacts on the transformers. Please note that the auxiliary cooling systems are additional systems that are not part of the transformer’s design, i.e., exclusive of cooling methods like AN, AF, ONAN, ONAF, etc. [115].
There have been mobile transformers that are used to mitigate the power grid impacts because of transformer failure in extreme temperatures [116]; this enabled to reduce the down time significantly. Another solution is to provide mobile auxiliary cooling services; for example, one substation in the eastern United States was facing 50,000 gallons of overheated transformer oil due to an overloaded transformer where the typical 208V/240V cooling application is not sufficient. In this case, portable trailer-mounted motor cooling was used to cool the transformer [117]. Another well-known substation solution is thermal imagery techniques to monitor the temperatures of the transformer and its adjacent equipment, which can also radiate heat increasing transformer’s temperature [110, 111, 112].
@@ VI-B Circuit Breaker
Circuit breakers are essential grid equipment whose role is to protect the system and isolate disturbances or faults in the power grid. Circuit breakers also carry power and have internal device heat during heavy loading conditions on hot days. The insulating medium in circuit breakers changes with the system voltage and has a role to play in reliable operation. High-voltage circuit breakers usually have an insulating medium: SF6 gas, nitrogen, or a mix of these gases. Vacuum circuit breakers are common high- and medium-voltage circuit breakers. Medium- and low-voltage circuit breakers are usually air-insulated circuit breakers. While reference [118] highlights that extreme cold conditions will degrade the performance of SF6 circuit breakers, reference [119] concludes that extreme heat and temperature rise of SF6 has little impact on the SF6 gas performance and its application in SF6 insulated power equipment. There are other studies in literature that have studied some gas mixtures like SF6–CO2 as insulating media in circuit breakers but do not discuss the performance under extreme temperatures [120].
@@
Impact of Extreme Temperature Operations: Medium-voltage and low-voltage circuit breakers that are mainly for industrial and residential protection are significantly affected by extreme heat and cold. Reference [121] provides details of de-rating and re-rating curves for low-voltage residential circuit breakers up to 63 A ratings. These clearly indicate that extreme heat and hot ambient air conditions will negatively impact the performance of the circuit breakers and can lead to tripping of the breakers much earlier than intended, leading to loss of load and often nuisance tripping. In aggregate, such behavior can cause significant impact on the load in the system leading, to unwanted cascaded actions on the distribution system level. Reference [122] discusses the reliability concerns of other low-voltage, high-power circuit breakers in extreme weather conditions. Evaluating the temperatures experienced during transportation and storage is an important consideration for circuit breakers, and needs to be documented to ensure the grid operators and asset owners understand the expected life and reliability vulnerability of the equipment that is being commissioned.
@@
Reference [123] is a report by NERC that discusses examples of circuit breaker failures due to extreme cold. These occurrences are important to consider in various regions of the country as the unstable polar vortex phenomenon is happening with increasing frequency and reaching regions of the country that normally did not expect extreme cold.
@@
Existing Methods to Measure the Impacts: Reference [124] Provides details of some methods for monitoring SF6 gas conditions in SF6 circuit breakers. Reference [125] provides a circuit breaker monitoring system that asset owners can use to monitor various operating conditions of the circuit breakers and can itself operate in temperatures ranging from -$40$°C to +65°C, which means that it might need additional thermal insulation to operate in extreme temperature exposures like during polar vortices or extreme hot days inside an enclosure. Other details specifications of this monitoring system can be found in reference [126].
@@
Existing Methods to Mitigate the Impacts: While extreme heat-related failures can be managed to some extent by derating that compensates for the temperature rise caused due to loading of the equipment, extreme cold conditions are difficult to managed for outdoor equipment. Reference [7] provides links to other learning and reported material associated with winter preparedness for circuit breakers and other critical grid equipment. Winterizing the equipment that depend on fluids is important, and often high-voltage circuit breakers have some media around the contacts to extinguish the arc and break the circuit [123].